evaluate

  1. "\int" cscnxdx
1
Expert's answer
2021-07-28T17:08:59-0400

We have



In = "\int cosec^n x \ dx"


In = "\int" cosecn-2 x cosec2 x dx


Now integrating the above expression by by using the formula of "\int" uv dx.



Considering cosecn-2 x = u and cosec2 x = v. We have,


In = cosecn-2 x "\int" cosec2 x dx - "\int" [ "\dfrac{d}{dx}" (cosecn-2 x) "\int" cosec2 x dx ] dx


In = - cosecn-2 x cot x - "\int" [ (n-2) cosecn-3 x cosec x cot2 x] dx


In = - cosecn-2 x cot x - "\int" [ (n-2) cosecn-2 x (cosec2 x - 1) ] dx


In = - cosecn-2 x cot x - (n - 2) "\int" [ cosecn x - cosecn-2 x ] dx ...............equation(1)


Noe from the above equation it can be seen that


In = "\int" cosecn x dx and In-2 = "\int" cosecn-2 x dx .................equation(2)


So on substituting the values from equation(2) to equation(1), we have



In = - cosecn-2 x cot x - (n - 2) [ In - In-2 ]


In (n - 1) = - cosecn-2 x cot x + (n - 2) In-2


In = "\dfrac{-1}{n-1}" cosecn-2 x cot x + "\dfrac{n-2}{n-1}" In-2